Using hooks law, calculate the wavenumber and wavelength of the IR radiation absorbed due to stretching vibration of the carbonyl group the masses of C and O are 19.92×10^-24 and 26.56×10^-24 respectively. The force constants for double bond is 10×10^5 gram per seconds sequare.


1
Expert's answer
2021-01-15T06:48:23-0500

k=2π/λ

According to hooke's law

F = kE

E = 26.56×10^-24 - 19.92×10^-24

=

k = F/E

k = 10×105 / 6.64g

k = 1.5 × 105(wavenumber)


λ =2π/k

= 2π/1.5 × 105

= 4.2× 10-5m(wavelength)